OpenStudy (anonymous):

When gas expands in a cylinder with radius 'r', the pressure at any given time is a function of the volume: P=P(V). The force exerted by the gas on the piston is the product of the pressure and the area: F=pi(r^2)P. Show that the work done by the gas when the volume expands from volume V1 to volume V2 is W= integral (V1 to V2) PdV. Medal to anyone who can help me show that this integral is correct!

OpenStudy (anonymous):

The definition of work is \[ W = \int \vec{F} \cdot d\vec{r} \] or, in your case, since everything is in one line, \[ W = \int F dz \] Where F is the pressure times the area and dz is the change in height of the piston. You can rewrite this in terms of pressure and volume.

OpenStudy (anonymous):

Because pressure is the force being exerted right?

OpenStudy (anonymous):

No, but it's related to the force... re-read what you wrote in the beginning.

OpenStudy (anonymous):

OH I totally understand now!! Thank you so much!!!